Given that the random "W" is binomial distribution with "n" trials and success probability "p" in each trial and "P(W=w)=h(w)",show that

(a)."h(w)\/h(w-1)=p(n-w+1)\/(1-p)w,w>0"

(b)."E[W(W-1)]=n(n-1)p"2

(c)."E(1\/W+1)=1-(1-p)""n+1" /(n+1)p

1
Expert's answer
2021-05-07T10:35:03-0400

Solution:

"P(W=w)=h(w)=^nC_wp^w(1-p)^{n-w}"

(a) "h(w-1)=^nC_{w-1}p^{w-1}(1-p)^{n-w+1}"

Now, "\dfrac{h(w)}{h(w-1)}=\dfrac{^nC_wp^w(1-p)^{n-w}}{^nC_{w-1}p^{w-1}(1-p)^{n-w+1}}"

"=\dfrac{\dfrac{n!}{w!(n-w)!}p}{\dfrac{n!}{(w-1)!(n-w+1)}(1-p)}"

"=\dfrac{(w-1)!(n-w+1)!p}{(w!)(n-w)!(1-p)}=\dfrac{(w-1)!(n-w+1)(n-w)!p}{w(w-1)!(n-w)!(1-p)}"

"=\dfrac{(n-w+1)p}{w(1-p)}"

(b) "E[W(W-1)]=E[W^2-W]=E(W^2)-E(W)" ...(i)

Now, "E(W)=np, Var(W)=E(W^2)-[E(W)]^2=np(1-p)"

So, "E(W^2)=[E(W)]^2+np(1-p)=n^2p^2+np(1-p)"

Putting these values in (i)

"E[W(W-1)]=E(W^2)-E(W)=n^2p^2+np(1-p)-np"

"=n^2p^2+np-np^2-np
\\=n^2p^2-np^2
\\=np^2(n-1)=n(n-1)p^2"

(c) "E[\dfrac1{W+1}]=\Sigma(\dfrac1{w+1})P(W=w)"

"=\Sigma_{w=0}^n(\dfrac1{w+1})\times ^nC_wp^w(1-p)^{n-w}"

"=\dfrac{1}{p(n+1)}\Sigma_{w=0}^n .^{n+1}C_{w+1}.p^{w-1}(1-p)^{n-w}"

"=\dfrac{1}{p(n+1)}.(1-(1-p)^{n+1})"